OpenStudy (anonymous):

Solve for x. 3(x + 2) + 4(x - 5) = 10 x = 36/7 x = 24/7 x = 16/7 x = 9/7

Parth (parthkohli):

-=> 3x + 6 + 4x - 20 = 10 => 7x - 14 = 10 => x = 24/7

OpenStudy (anonymous):

3(x + 2) + 4(x-5) = 10 3x + 6 + 4x - 20 = 10 7x - 14 = 10 7x = 24 x = 24/7